# Edge Graph Intelligence: Reciprocally Empowering Edge Networks with Graph Intelligence

### v1

@@ II-B General Workflow
II-B1 Preprocessing
The first step serves as an initialization to prepare data in a format aligned with the targeted GI model’s requirement. This can be, for example, reorganizing the adjacency matrix in a dense format or the compressed sparse row format and dropout some irrelevant elements from the feature vectors. Since the GI model is often known ahead of runtime, graph preprocessing is usually done offline.
II-B2 Sampling
With the preprocessed input graph, the GI model dives into iterations. The number of iterations required is exactly the number of layers the GI model possesses. Within each iteration, it first applies sampling on the input graph $\mathcal{G}$ to reduce the computational complexity of subsequent steps. Assuming the sampling function as $\theta^{(l)}$, this step can be formally written in:
@@ II-B2 Sampling
II-B3 Aggregation
Upon the sampled graph $\mathcal{G}^{\prime}$, the GI model performs neighbor aggregation where each vertex/link pulls feature vectors from its neighbors. Taking vertices aggregation as example, let $\varphi^{(l)}_{V}$ be the aggregation function of vertices, we have
@@ II-B3 Aggregation
II-B4 Update
The GI model then passes the aggregation $h^{(l)}_{v}$ through a neural network operator $\phi^{(l)}_{V}$ to update $v$’s representation vector:
@@ II-B4 Update
II-B5 Pooling
Pooling is also an optional step that aims at reducing the original graph to a smaller graph for lower computational complexity. It directly operates the sampled graph with updated representation vectors, pooling fields of graphs:
@@ II-B5 Pooling
II-B6 Readout
The above sampling, aggregation, update, and pooling steps will iterate until all model layers are processed, and thereafter generate embeddings of all vertices and links. To attain the desired results demanded by applications, these embeddings are obliged to the final readout step, which applies a pre-defined operator or model to transform embeddings to a global output, as explained in Sec. II-A5. Given a readout function $\psi$, the global output vector can be obtained by:
@@ II-B6 Readout
II-C Graph Learning Models
@@ II-C1 Recurrent Graph Neural Network (RecGNN)
RecGNNs are a pioneering architecture that builds the conceptual foundation in the field of graph representation learning [9]. They are primely proposed to learn node representations using recurrent neural architectures, integrating a recurrent hidden state and graph signal processing (GSP) to exploit the spatial structural information inherent in graph processes. As a genre of GI models dating back to the “pre-deep-learning” era, RecGNNs have inspired numerous subsequent research such as convolutional variants.
II-C2 Convolutional Graph Neural Network (ConvGNN)
@@
ConvGNNs extend convolution operations from grid data to graph data, aggregating the features of vertices and links with their neighbors to generate representations [10, 14]. Contrary to RecGNNs, successive graph convolutional layers are stacked in ConvGNNs to extract hierarchical patterns from subgraphs [16]. Among GNNs, ConvGNNs serve as a foundational component in constructing various advanced GI models [39].
II-C3 Graph Attention Network (GAT)
@@
GATs inherit spatial ConvGNNs by incorporating the attention mechanism into the aggregation functions, which effectively improves the capacity as well as the expressiveness of GI models [40]. The rationale behind this combination is to differentiate the contribution of vertices’ neighbors in a learning manner [41]. Based on this, many types of attention mechanisms are derived such as self-attention, gating attention, and semantic-level attention.
II-C4 Graph Autoencoder (GAE)
@@
GAEs are unsupervised frameworks that encode vertices, links, or graphs into a latent vector space and reconstruct graph data by decoding their encoded information [13]. GAEs are particularly useful for graph generation tasks, where a GAE model employs graph convolutional layers to compute embeddings for vertices and rebuild the graph adjacency matrix via decoders.
II-C5 Spatio-Temporal Graph Neural Network (STGNN)
@@
STGNNs analyze dynamic graphs from both the spatial and temporal dimensions [42, 43]. Each vertex and link in these graphs attaches a feature vector that describes their behaviors within a time window, and STGNNs aim to learn their patterns and predict their changes in the incoming time slots. To extract information from spatial-temporal dependencies, many sequential decision-making approaches are applied, e.g., Long Short-Term Memory (LSTM).
II-C6 Graph Transformer and Graph Foundation Model
@@
Graph transformers explore embracing transformer architecture to the graph domain in pursuit of improved graph modeling ability [44, 45, 46]. Typically, graph transformers incorporate GNN with transformers in dual ways: 1) design tailored positional embedding modules and graph-specific attention matrices, and 2) exploit GNNs as an auxiliary module by combining GNNs into transformer architectures [47]. Following this line, scaling small transformers to huge foundation models leads to Graph Foundation Models (GFMs) [48, 48] and Large Graph Models (LGMs) [49]. Motivated by the success of Large Language Models (LLMs), many researchers believe that utilizing pretraining techniques to resolve massive graph data can breed a comprehensive GI model, which possesses advanced abilities such as in-context graph understanding and versatile graph reasoning [50]. Nevertheless, GFMs and LGMs are still in a very early stage of their development and demands for further exploration.
II-D Learning Paradigms
@@ II-D1 Supervised Learning
Supervised learning is one of the most fundamental training paradigms in ML, which requires labeled data to guide the optimization of models [51, 52]. For GNNs, supervised learning enables to capture both local and global graph structures and the latent information of vertices and links. It iteratively optimizes model parameters by minimizing the difference between the model’s predictions and the labels, where the performance of the model is evaluated on a separate test set of labeled objects. Supervised GL have found success in various domains such as social network analysis, intelligent transportation, and recommendation systems. However, in many real-world scenarios where labels are unavailable or expensive, supervised learning is constrained and other semi-supervised or unsupervised learning paradigms are introduced.
II-D2 Transfer Learning (TL)
@@
TL for GI models builds upon the assumption that the learned representations and knowledge from a source graph can be effectively transferred and applied to a targeted graph [53, 54]. Based on this, it exploits knowledge distillation methods to extract the knowledge from one graph data to improve the learning performance on another different but related graph data. By transferring knowledge across graphs, the target graph benefits from the learned representations, enabling improved prediction performance for the targeted graph, even with limited labeled data.
II-D3 Contrastive Learning (CL)
@@
CL for GI models is a self-supervised learning paradigm that learns meaningful representations by contrasting positive and negative samples, where positive samples are pairs of vertices, links, or subgraphs that are similar or related, and negative ones are the contrary [55, 56, 57]. With these samples, contrastive learning maximizes the similarity between positive samples and minimizes the similarity between negative samples. Various techniques have been applied to enhance contrastive learning on graphs, such as graph augmentation and sample generation with random walks.
II-D4 Federated Graph Learning (FGL)
@@
FGL applies Federated Learning (FL) on graphs, which allows multiple clients to collaboratively train a GI model without sharing their local data [58, 59, 60]. In particular, each client trains the GI model using its local graph data and shares only the model updates, i.e., gradients or weights, with a central coordinator, and the coordinator aggregates the updates and sends a merged model update back to each client. Graph data are typically distributed across clients, with respect to structural segments, i.e., subgraphs, or feature segments. More details on FGL are discussed in Sec. V-A.
II-D5 Deep Graph Reinforcement Learning (DGRL)
@@
DGRL combines GI models with Deep Reinforcement Learning (DRL) techniques for interacting with graph environments [61, 62]. Typically, in DGRL, the GI model is responsible for processing the graph data, extracting features, and capturing the relationships between vertices and links. The DRL component then uses the embeddings computed by the GI model to learn a policy and make decisions, often by taking actions like vertex selection, link insertion/removal, or graph modification. Benefiting from the superior capability of representing graphs, DGRL has become a powerful tool for graph-based sequential decision-making tasks.
III Primer on Edge Networks
III-A Edge Networks and Edge Computing
@@
In the evolving landscape of distributed computing, the edge network emerges as a pivotal architecture, facilitating the transition from centralized cloud-based core networks to more decentralized edge networks [63, 64, 65]. Fig. 5 illustrates the edge network as a cohesive architecture that integrates the Edge and End layers, encompassing devices naturally situated closer to end users. Besides traditional network-style topology, edge networks can exhibit versatile organizations such as peer-to-peer connection (e.g., client-server mode) and star-like interactions (e.g., one-to-many subscription). Albeit, all these topologies can be unified by the graph abstraction. On the other hand, core networks and the Internet also comprise many graph structures and have introduced GI over their graphs. However, it should be emphasized that EGI is significantly different from GI on core networks given the unique characteristics of edge networks.
@@
Emerging from the advanced architecture of edge networks, edge computing stands as a new computing paradigm in contrast to cloud computing [8, 34]. Edge computing represents a shift in computational paradigm that brings data processing exponentially closer to the point of data collection and consumption, enabling low-latency and high-bandwidth communication essential for real-time applications. Edge computing, in comparison to cloud computing, presents numerous advantages, notably in its approach to in-situ data processing: (1) Lower Core Network Reliance: Edge computing lessens the reliance on unstable and delay-prone core networks by localizing data processing within edge networks, ensuring more consistent QoS even when wide-area network (WAN) connections are unpredictable. (2) Alleviated Core Network and Cloud Datacenter Stress: With the sheer volume of edge and end devices, routing all data through the core network to cloud data centers imposes substantial strain in terms of communication, computation, and storage. Edge computing mitigates this by retaining computations within the edge network, effectively utilizing the distributed, idle resources of edge and end devices. (3) Privacy and Security Enhancement: Localizing data within edge networks, minimizes the risk of unauthorized access and privacy breaches. Transferring these data to the cloud-based data centers owned by commercial companies inevitably raises users’ privacy concerns.
III-B Components of Edge Networks
@@
III-B1 Sensors and Micro Control Units
Sensors and Micro Control Units (MCUs) are fundamental components of edge networks. Sensors are devices that capture real-time data from the physical environment, such as temperature, motion, and gas composition. This type of devices are the most scattered IoT devices and thus located at the bottom level in Fig. 6, where several example sensors as shown. The DS18B20 [66], a digital temperature sensor, is known for its precision and ease of integration. The MPU-6050 [67] sensor tracks motion by combining a gyroscope and accelerometer, ideal for 3D movement analysis. Additionally, the MQ-2 [68] sensor detects combustible gases and smoke, illustrating the sensors’ diverse applications from home safety to industrial monitoring. MCUs are small computing devices that process data at the network’s edge. The Arduino Uno [69], a popular choice for DIY projects, excels in simple automation tasks. In contrast, the STM32F103C8T6 [70] is known for its robust performance in complex industrial applications. The ESP8266 [71], widely used in IoT devices, efficiently handles Wi-Fi connectivity. Positioned at the lowest level of the edge computing network’s hierarchy, these sensors and MCUs facilitate in-situ data collection and processing, significantly diminishing the necessity of relaying data to distant servers.
III-B2 Embedded and Mobile Devices
Embedded and Mobile Devices are advanced systems that integrate multiple sensors and MCUs into a single, often portable, framework. Unlike standalone sensors or MCUs that primarily capture or process data, these devices offer comprehensive functionality, including data processing, analysis, and user interaction. They excel in gathering, interpreting, and interacting with data, while seamlessly integrating with other devices and networks, showcasing advanced complexity and connectivity. Embedded devices are specialized computing units designed for specific functions within larger systems. The Raspberry Pi series [72] is a prime example, widely adopted for its versatility and accessibility in both educational settings and hobbyist projects. Advancing in performance, the NVIDIA Jetson series [73] steps in, equipped with GPUs to handle more resource-demanding tasks like AI and machine learning, making them suitable for complex computational needs. Transitioning to mobile devices, smartphones epitomize versatility and connectivity, with models like the iPhone [74] and Samsung Galaxy [75] offering advanced computing power and a wide array of features. Wearable technology, such as the Apple Watch [76] and Fitbit [77], further extends this mobility, providing health tracking and communication in portable, user-friendly designs.
@@ III-B2 Embedded and Mobile Devices
Figure 6: Spectrum of edge network can be classified into five categories: sensors and micro control units, embedded and mobile Devices, robotics and Vehicles, edge servers, and edge cloud.
III-B3 Robotics and Vehicles
Robotics and vehicles, distinct from embedded and mobile devices, usually excel in functionality, autonomy, and complexity. While embedded devices are integral to these systems, robotics and vehicles incorporate them into more complex frameworks, designed for higher autonomy and minimal human intervention. Robotics, for example, integrates sensors, embedded computers, and complex control algorithms to enable machines to autonomously interact with their environment. This autonomy is showcased in various applications, from industrial robots that perform precise and repetitive tasks, to service robots designed for dynamic environments like healthcare or customer support [78, 79], where they navigate, make decisions, and act with minimal human oversight. Autonomous vehicles, ranging from self-driving cars to unmanned aerial vehicles (UAVs) [80], are equipped with advanced navigation systems, real-time data processing, and sophisticated decision-making algorithms. These systems allow them to interpret and respond to diverse environmental conditions. For instance, autonomous cars [81] use a combination of LIDAR, radar, and cameras to understand their surroundings, while UAVs [82] might employ GPS and inertial navigation systems for precise positioning and path planning.
III-B4 Edge Servers
Edge servers act as micro-data centers to deliver services in a way that’s similar to using cloud servers but with a key difference: they are situated closer to the data source. This proximity not only ensures adherence to data localization laws but also significantly reduces data transfer latency. This is particularly beneficial for edge applications with stringent data privacy and real-time processing requirements. Edge servers, which typically exhibit modest performance in contrast to their cloud counterparts, are distinguished by their reduced power requirements and cost-effectiveness. A notable exemplification of edge servers includes Dell PowerEdge series [83] and Lenovo ThinkSystem series [84], both of which are industry standards for edge computing solutions. While edge servers often look like a typical server, many other form factors exist. Examples include a ruggedized laptop, a purpose-built appliance, or a robust mobile device featuring onboard intelligence, all of which can function as edge servers [85].
III-B5 Edge Cloud
@@
Edge cloud services extend cloud computing’s convenience to the network’s edge, with providers like AWS and Google offering solutions such as AWS Edge Service [86] and Google Distributed Cloud Edge [87]. Edge cloud are usually hosted by micro-data centers comprising edge servers that store, analyze, and process data faster than is possible using a connection to a cloud data center. By combining a distributed content delivery network (CDN) [88], the edge cloud significantly enhances services such as streaming for live broadcasting, cloud gaming, and virtual reality (VR) [89] experiences. These applications gain from the edge cloud’s reduction of data transmission latency, which concurrently amplifies local device storage and processing capabilities, yielding more efficient and agile digital experiences.
III-C Software Frameworks for Edge Networks
The environments in which edge networks operate are markedly more complex and heterogeneous compared to cloud data centers. This complexity is evident in multiple dimensions: (1) The connectivity landscape is fragmented with a plethora of communication protocols, such as MQTT, Zigbee, and Modbus [90], and is further complicated by the mobility and intermittence of many edge devices. (2) The diversity in operating systems is considerable, encompassing widely adopted systems such as Windows and Ubuntu, as well as specialized ones like Android, OpenWrt, HarmonyOS, and TinyOS [91], catering to various operational needs. (3) This heterogeneity is also evident in the platform architectures of edge nodes, which include a spectrum of architectures such as Arm, x86, and RISC-V [92], each contributing to the multifaceted nature of edge computing. (4) The development of IoT applications within this domain is characterized by a varied use of programming languages, with C/C++, Java, Python, and Go being predominant [93]. Moreover, edge computing environments are typically more resource-constrained than cloud computing in aspects such as computational power, memory, network bandwidth, and energy efficiency. These complexities and constraints pose significant challenges for software frameworks for edge networks, necessitating heterogeneity-aware and resource-optimized designs.
III-C1 Software Frameworks for Edge Computing

### v2

@@ III-A2 General Workflow
Preprocessing. The first step serves as an initialization to prepare data in a format aligned with the targeted GI model’s requirement. This can be, for example, reorganizing the adjacency matrix in a dense format or the compressed sparse row format and dropout some irrelevant elements from the feature vectors. Since the GI model is often known ahead of runtime, graph preprocessing is usually done offline.
Sampling. With the preprocessed input graph, the GI model dives into iterations. The number of iterations required is exactly the number of layers the GI model possesses. Within each iteration, it first applies sampling on the input graph $\mathcal{G}$ to reduce the computational complexity of subsequent steps. Assuming the sampling function as $\theta^{(l)}$, this step can be formally written in:
@@
Aggregation. Upon the sampled graph $\mathcal{G}^{\prime}$, the GI model performs neighbor aggregation where each vertex/link pulls feature vectors from its neighbors. Taking vertices aggregation as example, let $\varphi^{(l)}_{V}$ be the aggregation function of vertices, we have
@@
Update. The GI model then passes the aggregation $h^{(l)}_{v}$ through a neural network operator $\phi^{(l)}_{V}$ to update $v$’s representation vector:
@@
Pooling. Pooling is also an optional step that aims at reducing the original graph to a smaller graph for lower computational complexity. It directly operates the sampled graph with updated representation vectors, pooling fields of graphs:
@@
Readout. The above sampling, aggregation, update, and pooling steps will iterate until all model layers are processed, and thereafter generate embeddings of all vertices and links. To attain the desired results demanded by applications, these embeddings are obliged to the final readout step, which applies a pre-defined operator or model to transform embeddings to a global output, as explained in Sec. II. Given a readout function $\psi$, the global output vector can be obtained by:
@@
III-A3 Graph Learning Models
@@
Recurrent Graph Neural Network (RecGNN) is a pioneering architecture that builds the conceptual foundation in the field of graph representation learning [9]. They are primely proposed to learn node representations using recurrent neural architectures, integrating a recurrent hidden state and graph signal processing (GSP) to exploit the spatial structural information inherent in graph processes. As a genre of GI models dating back to the “pre-deep-learning” era, RecGNNs have inspired numerous subsequent research such as convolutional variants.
@@
Convolutional Graph Neural Network (ConvGNN) extend convolution operations from grid data to graph data, aggregating the features of vertices and links with their neighbors to generate representations [10, 14]. Contrary to RecGNNs, successive graph convolutional layers are stacked in ConvGNNs to extract hierarchical patterns from subgraphs [15]. Among GNNs, ConvGNNs serve as a foundational component in constructing various advanced GI models [48].
@@
Graph Attention Network (GAT) inherit spatial ConvGNNs by incorporating the attention mechanism into the aggregation functions, which effectively improves the capacity as well as the expressiveness of GI models [49]. The rationale behind this combination is to differentiate the contribution of vertices’ neighbors in a learning manner [50]. Based on this, many types of attention mechanisms are derived such as self-attention, gating attention, and semantic-level attention.
@@
Graph Autoencoder (GAE) are unsupervised frameworks that encode vertices, links, or graphs into a latent vector space and reconstruct graph data by decoding their encoded information [13]. GAEs are particularly useful for graph generation tasks, where a GAE model employs graph convolutional layers to compute embeddings for vertices and rebuild the graph adjacency matrix via decoders. Variational GAE goes beyond traditional GAE by transforming vertex embeddings as distribution, where each node’s embedding is represented by a mean and variance, allowing the model to capture uncertainties by sampling from these distributions [51].
@@
Spatio-Temporal Graph Neural Network (STGNN) analyze dynamic graphs from both the spatial and temporal dimensions [27, 52]. Each vertex and link in these graphs attaches a feature vector that describes their behaviors within a time window, and STGNNs aim to learn their patterns and predict their changes in the incoming time slots. To extract information from spatial-temporal dependencies, many sequential decision-making approaches are applied, e.g., Long Short-Term Memory (LSTM).
@@
Graph Transformer and Graph Foundation Model. Graph transformers explore embracing transformer architecture to the graph domain in pursuit of improved graph modeling ability [53, 54, 55]. Typically, graph transformers incorporate GNN with transformers in dual ways: 1) design tailored positional embedding modules and graph-specific attention matrices, and 2) exploit GNNs as an auxiliary module by combining GNNs into transformer architectures [56]. Following this line, scaling small transformers to huge foundation models leads to Graph Foundation Models (GFMs) [57, 57] and Large Graph Models (LGMs) [58]. Motivated by the success of Large Language Models (LLMs), many researchers believe that utilizing pretraining techniques to resolve massive graph data can breed a comprehensive GI model, which possesses advanced abilities such as in-context graph understanding and versatile graph reasoning [59]. Nevertheless, GFMs and LGMs are still in a very early stage of their development and demand for further exploration.
III-A4 Learning Paradigms
@@
Supervised Learning is one of the most fundamental training paradigms in ML, which requires labeled data to guide the optimization of models [60, 61]. For GNNs, supervised learning enables to capture both local and global graph structures and the latent information of vertices and links. It iteratively optimizes model parameters by minimizing the difference between the model’s predictions and the labels, where the performance of the model is evaluated on a separate test set of labeled objects. Supervised GL has found success in various domains such as social network analysis, intelligent transportation, and recommendation systems. However, in many real-world scenarios where labels are unavailable or expensive, supervised learning is constrained and other semi-supervised or unsupervised learning paradigms are introduced.
@@
Transfer Learning (TL) for GI models builds upon the assumption that the learned representations and knowledge from a source graph can be effectively transferred and applied to a targeted graph [62, 63]. Based on this, it exploits knowledge distillation methods to extract the knowledge from one graph data to improve the learning performance on another different but related graph data. By transferring knowledge across graphs, the target graph benefits from the learned representations, enabling improved prediction performance for the targeted graph, even with limited labeled data.
@@
Contrastive Learning (CL) for GI models is a self-supervised learning paradigm that learns meaningful representations by contrasting positive and negative samples, where positive samples are pairs of vertices, links, or subgraphs that are similar or related, and negative ones are the contrary [64, 65, 66]. With these samples, contrastive learning maximizes the similarity between positive samples and minimizes the similarity between negative samples. Various techniques have been applied to enhance contrastive learning on graphs, such as graph augmentation and sample generation with random walks.
Variational Learning extends the idea of variational inference to GNN, introducing a probabilistic framework to learn node or graph-level representations. This technique is mainly applied in Variational GAEs [51], where each node’s embedding is treated as a latent variable sampled from a learned distribution. The variational framework enables the model to capture uncertainties in graph data by learning distributions over node embeddings rather than fixed values. This allows for uncertainty quantification by modeling distributions over node embeddings, improving robustness to noisy data, and preventing overfitting through regularization. This leads to more expressive representations and enables generative tasks like predicting missing links or generating new graph structures [67].
@@
Federated Graph Learning (FGL). FGL applies Federated Learning (FL) on graphs, which allows multiple clients to collaboratively train a GI model without sharing their local data [68, 69, 70]. In particular, each client trains the GI model using its local graph data and shares only the model updates, i.e., gradients or weights, with a central coordinator, and the coordinator aggregates the updates and sends a merged model update back to each client. Graph data are typically distributed across clients, with respect to structural segments, i.e., subgraphs, or feature segments. More details on FGL are discussed in Sec. VI-A.
@@
Deep Graph Reinforcement Learning (DGRL). DGRL combines GI models with Deep Reinforcement Learning (DRL) techniques for interacting with graph environments [29, 30]. Typically, in DGRL, the GI model is responsible for processing the graph data, extracting features, and capturing the relationships between vertices and links. The DRL component then uses the embeddings computed by the GI model to learn a policy and make decisions, often by taking actions like vertex selection, link insertion/removal, or graph modification. Benefiting from the superior capability of representing graphs, DGRL has become a powerful tool for graph-based sequential decision-making tasks.
III-B Edge Networks and Edge computing
III-B1 Basic Concepts
@@
Edge networks have emerged as a pivotal architecture, facilitating the transition from centralized cloud-based core networks to the much more decentralized edge [71, 72, 73]. Fig. 4 illustrates the edge network as a cohesive architecture that integrates the Edge and End layers, encompassing devices naturally situated closer to end users. Besides traditional network-style topology, edge networks can exhibit versatile organizations such as peer-to-peer connection (e.g., client-server mode) and star-like interactions (e.g., one-to-many subscription).
@@
Emerging from the advanced architecture of edge networks, edge computing stands as a new computing paradigm in contrast to cloud computing [8, 44]. Edge computing represents a shift in computational paradigm that brings data processing exponentially closer to the point of data collection and consumption, enabling low-latency and high-bandwidth communication essential for real-time applications [74, 75]. With lower reliance on the core network, edge computing not only alleviates both the stress of centralized cloud and backhaul bandwidth usage but also enhances privacy preservation through localized data processing.
III-B2 Components of Edge Networks
@@
Sensors are devices that capture real-time data from the physical environment, such as image, temperature, motion, and gas composition. This type of device is the most scattered IoT and is thus located at the bottom level in Fig. 5, where several example sensors as shown.
Embedded and Mobile Devices are advanced systems that integrate multiple sensors and MCUs into a single, often portable, framework. Unlike standalone sensors or MCUs that primarily capture or process data, these devices offer comprehensive functionality, including data processing, analysis, and user interaction. Embedded devices are specialized computing units designed for specific functions within larger systems.
@@
Figure 5: The spectrum of edge networks can be classified into five categories: sensors and micro control units, embedded and mobile devices, robotics and Vehicles, edge servers, and edge cloud.
Robots and Vehicles usually excel in functionality, autonomy, and complexity. While embedded devices are integral to these systems, robotics and vehicles incorporate them into more complex frameworks, designed for higher autonomy and minimal human intervention.
Edge Servers act as micro data centers to deliver services in a way similar to using cloud servers but with a key difference: they are situated closer to the data source. This proximity not only ensures adherence to data localization laws but also significantly reduces data transfer latency.
@@
Edge Cloud services extend cloud computing’s convenience to the network’s edge, with providers like AWS and Google offering solutions such as AWS Edge Service [76] and Google Distributed Cloud Edge [77]. Edge clouds are usually hosted by micro-data centers comprising edge servers that store, analyze, and process data faster than is possible using a connection to a cloud data center.
III-B3 Software Frameworks for Edge Networks